# Dynamical Dark Energy Beyond Planck? Constraints from multiple CMB probes, DESI BAO and Type-Ia Supernovae

### v1

@@ II Methodology
Table 1: Ranges for the flat prior distributions imposed on the 8 free cosmological parameters. In the analyses involving ACT and SPT data (both on their own and in combination with WMAP), I employ a Gaussian prior $\tau=0.065\pm 0.015$ as detailed in Sec. II.
@@
  • DESI: BAO measurements released by DESI from observations of galaxies and quasars Adame et al. (2024b), as well as Lyman-$\alpha$ Adame et al. (2024c) tracers. These measurements are summarized in Tab. I of Ref. Adame et al. (2024a).
@@
  • ACT: CMB temperature, polarization and lensing spectra from the ACT collaboration, combined with a Gaussian prior on $\tau=0.065\pm 0.015$, as done in Ref. Aiola et al. (2020). Specifically, I use:
    • the ACT-DR4 likelihood Aiola et al. (2020) for observations of temperature and polarization (TT, TE, EE) spectra made by ACT from 2017 to 2021;
    • the ACT-DR6 lensing likelihood based on the gravitational lensing mass map covering 9400 deg2, reconstructed from CMB measurements by ACT from 2017 to 2021 Madhavacheril et al. (2024); Qu et al. (2024b). In my analysis, I include only the conservative range of lensing multipoles $76340<\ell<$ 763.
@@ III.1 Constraints from independent CMB experiments
Planck+DESI+PP $-0.827\pm 0.062$ ($-0.83^{+0.12}_{-0.12}$) $-0.71^{+0.30}_{-0.25}$ ($-0.71^{+0.50}_{-0.56}$) $-7.61$
Planck+DESI+DESy5 $-0.736\pm 0.068$ ($-0.74^{+0.14}_{-0.13}$) $-0.99^{+0.33}_{-0.28}$ ($-0.99^{+0.57}_{-0.64}$) $-14.50$
ACT+DESI+PP $-0.855\pm 0.064$ ($-0.86^{+0.13}_{-0.12}$) $-0.55^{+0.33}_{-0.27}$ ($-0.55^{+0.56}_{-0.60}$) $-4.38$
WMAP+ACT+DESI+PP $-0.850\pm 0.064$ ($-0.85^{+0.13}_{-0.12}$) $-0.55^{+0.31}_{-0.26}$ ($-0.55^{+0.56}_{-0.58}$) $-5.20$
ACT+DESI+DESy5 $-0.762\pm 0.070$ ($-0.76^{+0.14}_{-0.13}$) $-0.84^{+0.37}_{-0.29}$ ($-0.84^{+0.63}_{-0.67}$) $-12.70$
WMAP+ACT+DESI+DESy5 $-0.765^{+0.065}_{-0.072}$ ($-0.76^{+0.14}_{-0.13}$) $-0.81^{+0.35}_{-0.28}$ ($-0.81^{+0.61}_{-0.65}$) $-13.93$
SPT+DESI+PP $-0.94^{+0.12}_{-0.11}$ ($-0.94^{+0.22}_{-0.24}$) $-0.06\pm 0.53$ ($-0.1^{+1.0}_{-1.0}$) $-2.27$
WMAP+SPT+DESI+PP $-0.92\pm 0.11$ ($-0.92^{+0.22}_{-0.22}$) $-0.17^{+0.59}_{-0.45}$ ($-0.17^{+0.97}_{-1.1}$) $-0.49$
SPT+DESI+DESy5 $-0.79^{+0.13}_{-0.11}$ ($-0.79^{+0.22}_{-0.25}$) $-0.62\pm 0.58$ ($-0.6^{+1.1}_{-1.1}$) $-0.14$
WMAP+SPT+DESI+DESy5 $-0.80^{+0.12}_{-0.11}$ ($-0.80^{+0.22}_{-0.25}$) $-0.61\pm 0.57$ ($-0.6^{+1.1}_{-1.1}$) $-4.33$
@@
I begin by examining Planck-based constraints. These are visually displayed in the top row of Fig. 1, with the combination of Planck+DESI+PP shown in the top left panel and Planck+DESI+DESy5 in the top right panel. Unsurprisingly, I find a preference for DDE which is consistent with the findings widely documented in the literature. This preference is already evident in Planck+DESI+PP: I find $w_{0}=-0.827\pm 0.062$ and $w_{a}=-0.71^{+0.30}_{-0.25}$, both at 68% CL. As clearly seen from Fig. 1, the cosmological constant case falls well outside the 95% CL contours. However, in line with the narrative reviewed thus far, I confirm that the shift towards DDE is substantially increased when replacing PP with DESy5. Comparing the left and right top panels of Fig. 1, it is evident that the marginalized probability contours move towards more negative values of $w_{a}$ and towards quintessence-like values of $w_{0}$ that are further away from the cosmological constant. This shift is also evident in the numerical constraints: for Planck+DESI+DESy5 they read $w_{0}=-0.736\pm 0.068$ and $w_{a}=-0.99^{+0.33}_{-0.28}$ at 68% CL. An interesting aspect of my re-analysis of Planck CMB data is that I am using the Plik (PR3) lensing likelihood, whereas the baseline analysis performed by the DESI collaboration used the NPIPE-PR4 lensing map in conjunction with ACT-DR6. On the one hand, here I do not want to mix data from different CMB experiments because I aim to compare results from CMB probes that are as independent as possible (with the exception of Sec. III.3, where I briefly examine what happens when combining Planck, ACT, and SPT data together). On the other hand, my analysis shows that using Plik (PR3) lensing leads to essentially the same results documented in the literature. Therefore ACT-DR6 and the new NPIPE-PR4 lensing map have minimal to no impact on driving this preference for DDE.
@@
As the next step, I focus on the results based on the ACT temperature, polarization (ACT-DR4) and lensing (ACT-DR6) likelihoods. The marginalized contours for $w_{0}$ and $w_{a}$ are shown in the middle row of Fig. 1, with ACT(+WMAP)+DESI+PP in the left panel and ACT(+WMAP)+DESI+DESy5 in the right panel. In this case, different remarks are warranted. First and foremost, including WMAP’s large-scale temperature and polarization measurements does not significantly impact the constraints on the DE parameters. WMAP measurements provide information around the first acoustic peak and help break degeneracies between cosmological parameters (most notably the degeneracy between the spectral index and the baryon energy density, see also Refs. Aiola et al. (2020); Giarè et al. (2023a)). They mainly shift the constraints on these anomalous parameters in ACT back to values inferred by Planck, without substantially improving the constraining power on $w_{0}$ and $w_{a}$. More crucial is the role of SN data. The sample used for SN distance moduli measurements now plays a key role in assessing the preference for DDE. If we focus on the combination of ACT(+WMAP)+DESI+PP, it is clear from Fig. 1 that the cosmological constant case falls within the 95% CL contours. This is also confirmed by the numerical results, which for the most constraining combination, ACT+WMAP+DESI+PP, are $w_{0}=-0.85^{+0.13}_{-0.12}$ and $w_{a}=-0.55^{+0.56}_{-0.58}$ at 95% CL. This result is important because it demonstrates that CMB data significantly influence the shift towards DDE. When using ACT+WMAP+DESI+PP, the preference for DDE is notably weaker compared to Planck+DESI+PP, highlighting the crucial role that Planck data play in reinforcing this preference. On the other hand, considering ACT(+WMAP)+DESI+DESy5, I still find a substantial preference for a quintessence-like EoS that crosses into the phantom regime. This is evident both from Fig. 1 and from the numerical constraints: $w_{0}=-0.76^{+0.14}_{-0.13}$ and $w_{a}=-0.81^{+0.61}_{-0.65}$ at 95% CL. This result confirms the findings discussed in Ref. Efstathiou (2024) regarding the importance of DESy5 in driving this shift towards DDE. However, despite such a preference being quite significant, it appears reduced compared to Planck+DESI+DESy5. In this regard, an important final remark is that the preference for DDE is not reduced due to larger uncertainties. Comparing the marginalized probability contours obtained for ACT(+WMAP) (middle right panels of Fig. 1) with those for Planck (top panels of Fig. 1), we see that these combinations of datasets produce essentially the same precision in constraining the parameter space of the model, see also Tab 2. Overall, in ACT(+WMAP)-based results, a genuine shift towards $w_{a}\to 0$ and $w_{0}\to-1$ is responsible for the reduced preference.
@@
The final step I undertake is to analyze the SPT temperature and polarization spectra. In this case, I can already anticipate an overall loss of constraining power due to the significantly larger error bars characterizing the SPT data, as well as the absence of information on the lensing power spectrum. Unsurprisingly, this loss of constraining power (compared to both Planck and ACT-based constraints) is evident in the bottom panels of Fig. 1. The same panels also reveal that including or omitting WMAP data at large scales can lead to a noticeable improvement in the constraints on the parameter space, highlighting how the final constraining power results from a comparable contribution from both WMAP and SPT. The larger uncertainties obtained in inferring cosmological parameters certainly represent a concern as they introduce significant correlations between the dark energy parameters $w_{0}$-$w_{a}$ and other crucial parameters characterizing the late time dynamics such as $H_{0}$ and $\Omega_{m}$. This is clearly seen in Fig. 4 and Fig. 5 in Appendix A. However, despite these caveats, it is important to stress that in this case no preference for DDE is found. The results for SPT(+WMAP)+DESI+PP reads $w_{0}=-0.92\pm 0.11$ and $w_{a}=-0.17^{+0.59}_{-0.45}$ at 68% CL. For SPT(+WMAP)+DESI+DESy5 I get $w_{0}=-0.80^{+0.12}_{-0.11}$ and $w_{a}=-0.61\pm 0.57$ at 68% CL. Both agree with $\Lambda$CDM within $\sim 2\sigma$. This is also evident in the bottom panels of Fig. 1: whether combining SPT(+WMAP)+DESI+PP or SPT(+WMAP)+DESI+DESy5, the resulting marginalized contours in the $w_{0}$-$w_{a}$ plane remain consistent with the cosmological constant, within or at the border of the 95% CL contours. Additionally, the choice of SN sample does not appear to have a significant impact. Although one might argue that any potential preference for DDE is buried in the larger SPT uncertainties, this result provides further concrete evidence that CMB data play a relevant role in reducing or increasing the statistical preference for DDE.
@@
Overall, my analysis reveals a trend in CMB experiments other than Planck to significantly reduce the preference for DDE. This implies that the preference is partially driven by Planck CMB observations. This trend is also evident when examining the $\Delta\chi^{2}=\chi^{2}_{\text{CPL}}-\chi^{2}_{\Lambda\text{CDM}}$ (last column of Tab 2) which represents the difference between the best-fit $\chi^{2}$ obtained using a CPL parameterization and within the standard $\Lambda$CDM model of cosmology. For combinations of datasets involving Planck, I observe a significant improvement in the fit over $\Lambda$CDM, with $\Delta\chi^{2}\sim-7.61$ ($\Delta\chi^{2}\sim-14.50$) for Planck+DESI+PP (Planck+DESI+DESy5). However, this improvement is reduced to $\Delta\chi^{2}\sim-4.38$ ($-5.20$) for ACT(+WMAP)+DESI+PP and becomes significant again with $\Delta\chi^{2}\sim-12.70$ ($-13.93$) for ACT(+WMAP)+DESI+DESy5. For SPT, the improvement is very modest, with values ranging from $\Delta\chi^{2}\sim-0.5$ to $-2$ for ACT(+WMAP)+DESI+PP and $\Delta\chi^{2}\sim-0.1$ to $-4$ for ACT(+WMAP)+DESI+DESy5. Thus, across all results based on different CMB experiments, the $\Delta\chi^{2}$ follows the same trend as the preference for DDE: the improvement in $\Delta\chi^{2}$ over $\Lambda$CDM is reduced when the preference for DDE diminishes. This suggests that the shift in the constraints on the parameter space from different CMB data is not merely due to a volume effect but is driven by differences in the fit, further underscoring the significant role that CMB data play in assessing the strength of the preference. Lastly, my analysis reaffirms the crucial role of DESy5 supernovae, as already discussed in the literature. Notably, this is the only dataset that reproduces evidence for DDE in the ACT-based constraints.
@@ III.2 Interpreting the Planck preference for DDE
A good starting point is to note that the most significant impact on the CMB angular power spectra resulting from the DE dynamics is primarily due to changes in the amplitude of the ISW plateau at very large angular scales. This amplitude is mainly controlled by primordial inflationary parameters (i.e., $A_{s}$ and $n_{s}$) and contributions from the late-time ISW effect. The latter is sensitive to DE dynamics, as a phantom (quintessence) DE component can oppose (favor) the decay of gravitational potentials that source the late ISW effect. Additionally, as highlighted several times in the literature, temperature and polarization data at large angular scales ($\ell\leq 30$) are crucial for understanding several mild anomalies observed in Planck’s CMB measurements Giarè et al. (2024c). This includes a latent preference for a phantom-like dark energy component Aghanim et al. (2020a); Escamilla et al. (2024b), which is closely related to large-scale temperature and E-mode polarization measurements Escamilla et al. (2024b); Giarè et al. (2024c). Given these premises, I have a priori reason to hypothesize that part of Planck’s tendency to strengthen the preference for DDE could result from large-scale temperature and polarization measurements as well. To corroborate this hypothesis, I performed the following tests: I compare the results obtained within the two baseline datasets involving Planck (Planck+DESI+PP and Planck+DESI+DESy5) with the results obtained for the same combination of data, removing information at low multipoles (i.e., large angular scales). In particular, I consider the following reduced likelihoods for the Planck spectra:
@@
In the top panel of the figure, for reference, I plot the results obtained within the baseline combination of data (Planck+DESI+PP and Planck+DESI+DESy5) where all information at large scales is retained. In the second panel from the top, I show the two-dimensional probability contours obtained for Planck(no low-T)+DESI+PP and Planck(no low-T)+DESI+DESy5 – i.e., removing the $-\ell$ commander likelihood for the TT spectrum. As seen in the figure, the constraints slightly shift toward the cosmological constant value for both datasets, but the effect is very modest. I obtain $w_{0}=-0.83^{+0.13}_{-0.12}$ and $w_{a}=-0.72^{+0.51}_{-0.58}$ for Planck(no low-T)+DESI+PP, and $w_{0}=-0.73^{+0.14}_{-0.13}$ and $w_{a}=-1.00^{+0.57}_{-0.65}$ for Planck(no low-T)+DESI+DESy5 (all at 95% CL), underscoring that the preference for DDE reduces, but remains solid. A more significant effect is seen in the third panel from the top. In this case, I am removing only the $-\ell$ SimAll likelihood for the EE spectrum. As I and other collaborators argued in Refs. Escamilla et al. (2024b); Giarè et al. (2024c), large-scale E-mode polarization measurements are of crucial importance for the constraints on the DE EoS. In this case, remarkably, the cosmological constant value falls within the 95% CL contours for Planck(no low-E)+DESI+PP, where I obtain $w_{0}=-0.85^{+0.13}_{-0.12}$ and $w_{a}=-0.56^{+0.55}_{-0.58}$ at 95% CL. I would like to underline the significant shift in the central value of $w_{a}$. A preference for a present-day quintessence-like EoS that turned phantom in the past remains substantial for Planck(no low-E)+DESI+DESy5: $w_{0}=-0.76^{+0.14}_{-0.13}$ and $w_{a}=-0.84^{+0.63}_{-0.66}$ at 95% CL. However, this preference is clearly reduced compared to the baseline case, again due to a shift in the value of $w_{a}$. When I remove any Planck temperature and polarization measurements at $\ell\lesssim 30$, the preference for DDE is further reduced. Referring to the bottom panel of Fig. 2, we see that the cosmological constant value falls within the 95% CL regions for Planck(no low-L)+DESI+PP, where I obtain $w_{0}=-0.85\pm 0.13$ and $w_{a}=-0.55^{+0.56}_{-0.61}$ at 95% CL. Conversely, the combination Planck(no low-L)+DESI+DESy5 still shows a significant preference for a dynamical EoS, yielding $w_{0}=-0.76^{+0.14}_{-0.13}$ and $w_{a}=-0.84^{+0.63}_{-0.67}$ at 95% CL. That said, when removing $-\ell$ data, the shift in the value of $w_{a}$ toward less negative values remains.
Given these results, there is solid ground to conclude that Planck temperature and polarization measurements at large angular scales $\ell\lesssim 30$ (especially E-mode polarization measurements) further strengthen the preference for DDE observed in DESI BAO and Type-Ia SN. Indeed, by removing these, the preference is reduced, and Planck(no low-L)-based constraints behave similarly to ACT-based constraints.

### v2

@@ II Methodology
Table 1: Ranges for the flat prior distributions imposed on the 8 free cosmological parameters. In the analyses involving ACT and SPT data (both on their own and in combination with WMAP), I employ a Gaussian prior $\tau=0.065\pm 0.015$ as detailed in Sec. II. I examine the impact of this prior on these combinations in Appendix B.
@@
  • DESI: BAO measurements released by DESI based on its second year of observations (DR2) of galaxies and quasars Andrade et al. (2025), as well as Lyman-$\alpha$ Abdul Karim et al. (2025b) tracers. These measurements are summarized in Tab. IV of Ref. Abdul Karim et al. (2025a). [^4] In the analysis, I included the full correlation matrix among different distance measurements using the publicly available likelihoods released for the sampler cobaya.
@@
  • ACT: latest CMB temperature, polarization and lensing spectra from the ACT collaboration, combined with a Gaussian prior on $\tau=0.065\pm 0.015$, as done in Ref. Aiola et al. (2020). Specifically, I use:
    • the ACT-DR6 likelihood Naess et al. (2025); Louis et al. (2025); Calabrese et al. (2025) for observations of temperature and polarization (TT, TE, EE) spectra measured from the Data Release 6 maps made from ACT data. These cover 19,000 deg2 of sky in bands centered at 98, 150, and 220 GHz, with white noise levels three times lower than Planck in polarization Louis et al. (2025);
    • the ACT-DR6 lensing likelihood based on the gravitational lensing mass map covering 9400 deg2, reconstructed from CMB measurements by ACT from 2017 to 2021 Madhavacheril et al. (2024); Qu et al. (2024). In my analysis, I include only the conservative range of lensing multipoles $76340<\ell<$ 763.
@@ III.1 Constraints from independent CMB experiments
Planck+DESI+PP $-0.846\pm 0.054$ ($-0.85^{+0.11}_{-0.11}$) $-0.56^{+0.22}_{-0.19}$ ($-0.56^{+0.39}_{-0.42}$) $-7.79$
Planck+DESI+DESy5 $-0.760\pm 0.057$ ($-0.76^{+0.11}_{-0.11}$) $-0.80^{+0.24}_{-0.21}$ ($-0.80^{+0.42}_{-0.46}$) $-19.20$
ACT+DESI+PP $-0.851\pm 0.056$ ($-0.85^{+0.11}_{-0.11}$) $-0.52^{+0.24}_{-0.20}$ ($-0.52^{+0.41}_{-0.46}$) $-6.14$
WMAP+ACT+DESI+PP $-0.859\pm 0.055$ ($-0.86^{+0.11}_{-0.11}$) $-0.47^{+0.22}_{-0.20}$ ($-0.47^{+0.40}_{-0.44}$) $-5.81$
ACT+DESI+DESy5 $-0.765\pm 0.058$ ($-0.77^{+0.12}_{-0.11}$) $-0.77^{+0.26}_{-0.23}$ ($-0.77^{+0.46}_{-0.49}$) $-16.77$
WMAP+ACT+DESI+DESy5 $-0.774\pm 0.057$ ($-0.77^{+0.11}_{-0.11}$) $-0.71\pm 0.23$ ($-0.71^{+0.44}_{-0.47}$) $-15.96$
SPT+DESI+PP $-0.888\pm 0.057$ ($-0.89^{+0.11}_{-0.11}$) $-0.27^{+0.26}_{-0.23}$ ($-0.27^{+0.46}_{-0.50}$) $-3.58$
WMAP+SPT+DESI+PP $-0.882\pm 0.056$ ($-0.88^{+0.11}_{-0.11}$) $-0.29^{+0.25}_{-0.22}$ ($-0.29^{+0.43}_{-0.48}$) $-4.33$
SPT+DESI+DESy5 $-0.799\pm 0.061$ ($-0.80^{+0.12}_{-0.12}$) $-0.55^{+0.28}_{-0.25}$ ($-0.55^{+0.50}_{-0.54}$) $-14.73$
WMAP+SPT+DESI+DESy5 $-0.798\pm 0.061$ ($-0.80^{+0.12}_{-0.11}$) $-0.55^{+0.28}_{-0.24}$ ($-0.55^{+0.48}_{-0.54}$) $-14.67$
@@
[…]
@@
As the next step, I focus on the results based on the ACT temperature, polarization and lensing (ACT-DR6) likelihoods. The marginalized contours for $w_{0}$ and $w_{a}$ are shown in the middle row of Fig. 1, with ACT(+WMAP)+DESI+PP in the left panel and ACT(+WMAP)+DESI+DESy5 in the right panel. In this case, different remarks are warranted. First and foremost, including WMAP’s large-scale temperature and polarization measurements does not significantly impact the constraints on the DE parameters. WMAP measurements provide information around the first acoustic peak and help break degeneracies between cosmological parameters (most notably the degeneracy between the spectral index and the baryon energy density, see also Refs. Aiola et al. (2020); Giarè et al. (2023a)). They mainly shift the constraints on these anomalous parameters in ACT back to values inferred by Planck, without substantially improving the constraining power on $w_{0}$ and $w_{a}$. More crucial is the role of SN data. The sample used for SN distance moduli measurements now plays a key role in assessing the preference for DDE. If we focus on the combination of ACT(+WMAP)+DESI+PP, it is clear from Fig. 1 that the cosmological constant case falls close to the 95% CL contours. This is also confirmed by the numerical results, which for the most constraining combination, ACT+WMAP+DESI+PP, are $w_{0}=-0.859\pm 0.055$ and $w_{a}=-0.47^{+0.22}_{-0.20}$ at 68% CL. This result is important because it demonstrates that CMB data significantly influence the shift towards DDE. When using ACT+WMAP+DESI+PP, the preference for DDE is notably weaker compared to Planck+DESI+PP, highlighting the crucial role that Planck data play in reinforcing this preference. On the other hand, considering ACT(+WMAP)+DESI+DESy5, I still find a substantial preference for a quintessence-like EoS that crosses into the phantom regime. This is evident both from Fig. 1 and from the numerical constraints: $w_{0}=-0.774\pm 0.057$ and $w_{a}=-0.71\pm 0.23$ at 68% CL. This result confirms the findings discussed in Ref. Efstathiou (2025) regarding the importance of DESy5 in driving this shift towards DDE. However, despite such a preference being quite significant, it appears reduced compared to Planck+DESI+DESy5. In this regard, an important final remark is that the preference for DDE is not reduced due to larger uncertainties. Comparing the marginalized probability contours obtained for ACT(+WMAP) (middle right panels of Fig. 1) with those for Planck (top panels of Fig. 1), we see that these combinations of datasets produce essentially the same precision in constraining the parameter space of the model, see also Tab 2. Overall, in ACT(+WMAP)-based results, a genuine shift towards $w_{a}\to 0$ and $w_{0}\to-1$ is responsible for the reduced preference.
@@
The final step I undertake is to analyze the SPT temperature and polarization spectra. In this case, I can already anticipate an overall loss of constraining power due to the significantly larger error bars characterizing the SPT data, as well as the absence of information on the lensing power spectrum. The larger uncertainties obtained in inferring cosmological parameters certainly represent a concern, as they introduce significant correlations between the dark energy parameters $w_{0}$$w_{a}$ and other crucial parameters characterizing the late-time dynamics such as $H_{0}$ and $\Omega_{m}$. This is clearly seen in Fig.4 and Fig.5 in Appendix A. However, despite these caveats, it is important to stress that in this case no strong preference for DDE is found in combinations involving Pantheon-plus SN. The results for SPT(+WMAP)+DESI+PP read $w_{0}=-0.882\pm 0.056$ and $w_{a}=-0.29^{+0.25}_{-0.22}$ at 68% CL – in agreement with $\Lambda$CDM within $\sim 2\sigma$. This is also evident in the left bottom panels of Fig. 1. Instead, for SPT(+WMAP)+DESI+DESy5 I get $w_{0}=-0.798\pm 0.061$ and $w_{a}=-0.55^{+0.28}_{-0.24}$ at 68% CL. Overall, whether combining SPT(+WMAP)+DESI+PP or SPT(+WMAP)+DESI+DESy5, the resulting marginalized contours in the $w_{0}$$w_{a}$ plane shift significantly closer to the cosmological constant value, even remaining within or at the border of the 95% CL contours. This outcome reinforces the important role that CMB measurements play in modulating the statistical evidence for DDE.
@@
Overall, my analysis reveals a trend in CMB experiments other than Planck to significantly reduce the preference for DDE. This implies that the preference is partially driven by Planck CMB observations. This trend is also evident when examining the $\Delta\chi^{2}=\chi^{2}_{\text{CPL}}-\chi^{2}_{\Lambda\text{CDM}}$ (last column of Tab 2) which represents the difference between the best-fit $\chi^{2}$ obtained using a CPL parameterization and within the standard $\Lambda$CDM model of cosmology. For combinations of datasets involving Planck, I observe a significant improvement in the fit over $\Lambda$CDM, with $\Delta\chi^{2}\sim-7.79$ ($\Delta\chi^{2}\sim-19.20$) for Planck+DESI+PP (Planck+DESI+DESy5). However, this improvement is reduced to $\Delta\chi^{2}\sim-6.14$ ($-5.81$) for ACT(+WMAP)+DESI+PP and becomes significant again with $\Delta\chi^{2}\sim-16.77$ ($-15.96$) for ACT(+WMAP)+DESI+DESy5. Similarly, for SPT, the improvement is very modest, with values ranging from $\Delta\chi^{2}\sim-3.58$ to $-4.33$ for SPT(+WMAP)+DESI+PP and deceased to $\Delta\chi^{2}\sim-14.7$ for SPT(+WMAP)+DESI+DESy5. Thus, across all results based on different CMB experiments, the $\Delta\chi^{2}$ follows the same trend as the preference for DDE: the improvement in $\Delta\chi^{2}$ over $\Lambda$CDM is reduced when the preference for DDE diminishes. This suggests that the shift in the constraints on the parameter space from different CMB data is not merely due to a volume effect but is driven by differences in the fit, further underscoring the significant role that CMB data play in assessing the strength of the preference. Lastly, my analysis reaffirms the crucial role of DESy5 supernovae, as already discussed in the literature. Notably, this is the only dataset that reproduces evidence for DDE in the ACT-based constraints.
@@ III.2 Interpreting the Planck preference for DDE
A good starting point is to note that the most significant impact on the CMB angular power spectra resulting from the DE dynamics is primarily due to changes in the amplitude of the ISW plateau at very large angular scales. This amplitude is mainly controlled by primordial inflationary parameters (i.e., $A_{s}$ and $n_{s}$) and contributions from the late-time ISW effect. The latter is sensitive to DE dynamics, as a phantom (quintessence) DE component can oppose (favor) the decay of gravitational potentials that source the late ISW effect. Additionally, as highlighted several times in the literature, temperature and polarization data at large angular scales ($\ell\leq 30$) are crucial for understanding several mild anomalies observed in Planck’s CMB measurements Giarè et al. (2024c). This includes a latent preference for a phantom-like dark energy component Aghanim et al. (2020a); Escamilla et al. (2024b), which is closely related to large-scale temperature and E-mode polarization measurements Escamilla et al. (2024b); Giarè et al. (2024c), see also Appendix B for further discussions. Given these premises, I have a priori reason to hypothesize that part of Planck’s tendency to strengthen the preference for DDE could result from large-scale temperature and polarization measurements as well. To corroborate this hypothesis, I performed the following tests: I compare the results obtained within the two baseline datasets involving Planck (Planck+DESI+PP and Planck+DESI+DESy5) with the results obtained for the same combination of data, removing information at low multipoles (i.e., large angular scales). In particular, I consider the following reduced likelihoods for the Planck spectra:
@@
In the top panel of the figure, for reference, I plot the results obtained within the baseline combination of data (Planck+DESI+PP and Planck+DESI+DESy5) where all information at large scales is retained. In the second panel from the top, I show the two-dimensional probability contours obtained for Planck(no low-T)+DESI+PP and Planck(no low-T)+DESI+DESy5 – i.e., removing the $-\ell$ commander likelihood for the TT spectrum. As seen in the figure, the constraints slightly shift toward the cosmological constant value for both datasets, but the effect is very modest. I obtain $w_{0}=-0.844\pm 0.054$ and $w_{a}=-0.58^{+0.22}_{-0.19}$ for Planck(no low-T)+DESI+PP, and $w_{0}=-0.759\pm 0.056$ and $w_{a}=-0.81^{+0.23}_{-0.21}$ for Planck(no low-T)+DESI+DESy5 (all at 68% CL), underscoring that the preference for DDE reduces, but remains solid. A more significant effect is seen in the third panel from the top. In this case, I am removing only the $-\ell$ SimAll likelihood for the EE spectrum. As I and other collaborators argued in Refs. Escamilla et al. (2024b); Giarè et al. (2024c), large-scale E-mode polarization measurements are of crucial importance for the constraints on the DE EoS. In this case, remarkably, the cosmological constant value falls within the 95% CL contours for Planck(no low-E)+DESI+PP, where I obtain $w_{0}=-0.867\pm 0.055$ and $w_{a}=-0.41^{+0.24}_{-0.20}$ at 68% CL. I would like to underline the significant shift in the central value of $w_{a}$. A preference for a present-day quintessence-like EoS that turned phantom in the past remains substantial for Planck(no low-E)+DESI+DESy5: $w_{0}=-0.783\pm 0.058$ and $w_{a}=-0.66^{+0.26}_{-0.22}$ at 68% CL. However, this preference is clearly reduced compared to the baseline case, again due to a shift in the value of $w_{a}$. When I remove any Planck temperature and polarization measurements at $\ell\lesssim 30$, the preference for DDE is further reduced. Referring to the bottom panel of Fig. 2, we see that the cosmological constant value falls within the 95% CL regions for Planck(no low-L)+DESI+PP, where I obtain $w_{0}=-0.868\pm 0.056$ and $w_{a}=-0.41^{+0.23}_{-0.21}$ at 68% CL. Conversely, the combination Planck(no low-L)+DESI+DESy5 still shows a significant preference for a dynamical EoS, yielding $w_{0}=-0.782\pm 0.058$ and $w_{a}=-0.67^{+0.25}_{-0.22}$ at 68% CL. That said, when removing $-\ell$ data, the shift in the value of $w_{a}$ toward less negative values remains.
Given these results, there is solid ground to conclude that Planck temperature and polarization measurements at large angular scales $\ell\lesssim 30$ (especially E-mode polarization measurements, see Appendix B.) further strengthen the preference for DDE observed in DESI BAO and Type-Ia SN. Indeed, by removing these, the preference is reduced, and Planck(no low-L)-based constraints behave similarly to ACT-based constraints.